Determine the value of c that makes the function f (x, y)= c(x + y) a joint probability density function over the range 0 < x < 3 and 0 < y < x. Determine the following:
a)P(X<1,Y<2)
b) P(1<x<2)
(c)P(Y>1) (e)E(X)
(d) P(X<2,Y<2)

(f) E(Y) (g) Marginal probability distribution of X

(h) Conditional probability distribution of Y given X = 1
(i) E (Y | X = 1) (j) P (Y > 2 | X = 1)
(k) Conditional probability distribution of X given Y = 2
1
Expert's answer
2020-05-25T21:17:21-0400
"f(x,y)=c(x+y),0<x<3,0<y<x"

"\displaystyle\int_{0}^3\displaystyle\int_{0}^xf(x,y)dydx=\displaystyle\int_{0}^3\displaystyle\int_{0}^xc(x+y)dydx="

"=c\displaystyle\int_{0}^3\big[xy+{y^2 \over 2}\big]\begin{matrix}
 x \\
 0
\end{matrix}dx={3c \over 2}\displaystyle\int_{0}^3x^2dx={c \over 2}[x^3]\begin{matrix}
 3 \\
 0
\end{matrix}={27c \over 2}=1"

"c={2 \over 27}"

"f(x,y)={2 \over 27}(x+y),0<x<3,0<y<x"

a)


"P(X<1,Y<2)=\displaystyle\int_{0}^1\displaystyle\int_{0}^2{2 \over 27}(x+y)dydx="

"={2 \over 27}\displaystyle\int_{0}^1\big[xy+{y^2 \over 2}\big]\begin{matrix}
 2 \\
 0
\end{matrix}dx={2 \over 27}\displaystyle\int_{0}^1(2x+2)dx="

"={2 \over 27}\big[x^2+2x\big]\begin{matrix}
 1 \\
 0
\end{matrix}={2 \over 9}"

b)


"P(1<X<2)=P(1<X<2,0<Y<x)="

"=\displaystyle\int_{1}^2\displaystyle\int_{0}^x{2 \over 27}(x+y)dydx={2 \over 27}\displaystyle\int_{1}^2\big[xy+{y^2 \over 2}\big]\begin{matrix}
 x \\
 0
\end{matrix}dx="

"={1 \over 9}\displaystyle\int_{1}^2x^2dx={1 \over 27}\big[x^3\big]\begin{matrix}
 2 \\
 1
\end{matrix}={7 \over 27}"

(c)


"P(Y>1)=P(0<X<3,1<Y<x)="

"=\displaystyle\int_{0}^3\displaystyle\int_{1}^x{2 \over 27}(x+y)dydx={2 \over 27}\displaystyle\int_{0}^3\big[xy+{y^2 \over 2}\big]\begin{matrix}
 x \\
 1
\end{matrix}dx="

"={2 \over 27}\displaystyle\int_{0}^3({3 \over 2}x^2-x-{1 \over 2})dx={1 \over 27}\big[x^3-x^2-x\big]\begin{matrix}
 3 \\
 0
\end{matrix}="

"={5 \over 9}"

(d)


"P(X<2,Y<2)=P(0<X<2,0<Y<2)="

"=\displaystyle\int_{0}^2\displaystyle\int_{0}^2{2 \over 27}(x+y)dydx={2 \over 27}\displaystyle\int_{0}^2\big[xy+{y^2 \over 2}\big]\begin{matrix}
 2 \\
 0
\end{matrix}dx="

"={2 \over 27}\displaystyle\int_{0}^2(2x+2)dx={2 \over 27}\big[x^2+2x\big]\begin{matrix}
 2 \\
 0
\end{matrix}={16 \over 27}"

(e)


"E(X)=\displaystyle\int_{0}^3\displaystyle\int_{0}^x{2 \over 27}x(x+y)dydx="

"={2 \over 27}\displaystyle\int_{0}^3\big[x^2y+{xy^2 \over 2}\big]\begin{matrix}
 x \\
 0
\end{matrix}dx={1 \over 9}\displaystyle\int_{0}^3x^3dx="

"={1 \over 9}\big[{x^4 \over 4}\big]\begin{matrix}
 3 \\
 0
\end{matrix}={9 \over 4}"

(f)


"E(Y)=\displaystyle\int_{0}^3\displaystyle\int_{0}^x{2 \over 27}y(x+y)dydx="

"={2 \over 27}\displaystyle\int_{0}^3\big[{xy^2 \over 2}+{y^3 \over 3}\big]\begin{matrix}
 x \\
 0
\end{matrix}dx={5 \over 81}\displaystyle\int_{0}^3x^3dx="

"={5 \over 81}\big[{x^4 \over 4}\big]\begin{matrix}
 3 \\
 0
\end{matrix}={5 \over 4}"

(g) Marginal probability distribution of X 


"f_X(x)=\displaystyle\int_{0}^x{2 \over 27}(x+y)dy={2 \over 27}\big[xy+{y^2 \over 2}\big]\begin{matrix}
 x \\
 0
\end{matrix}={1 \over 9}x^2"

(h) Conditional probability distribution of Y given X = 1 


"f_{Y|X=1}(y)={f_{XY}(1,y) \over f_X(1)}={{2 \over 27}(1+y) \over{1 \over 9}(1)^2}={2 \over 3}(1+y)"

(i)


"E(Y|X=1)=\displaystyle\int_{1}^3yf_{Y|x=1}(y)dy=\displaystyle\int_{1}^3{2 \over 3}y(1+y)dy="

"={2 \over 3}\big[{y^2 \over 2}+{y^3 \over 3}\big]\begin{matrix}
 3 \\
 1
\end{matrix}={76 \over 9}"

(j)


"P(Y>2|X=1)=0"

(k) Conditional probability distribution of X given Y = 2

Marginal probability distribution of Y 


"f_Y(y)=\displaystyle\int_{0}^3{2 \over 27}(x+y)dx={2 \over 27}\big[xy+{x^2 \over 2}\big]\begin{matrix}
 3 \\
 0
\end{matrix}="

"={2 \over 27}(3y+{9 \over 2})"

"f_{X|Y=2}(x)={f_{XY}(x,2) \over f_Y(2)}={{2 \over 27}(x+2) \over{2 \over 3}(1+2)}={1 \over 27}(x+2)"